OpenStudy (anonymous):

please can someone explain what open balls, closed balls and spares are in a metric space

OpenStudy (zzr0ck3r):

The best way to think about it, so that it makes sense with the name, is with the standard metric on \(\mathbb{R}^2\). Give me a radius \(\delta\) and some point \(a\) and the open ball about \(a\) is all the points within \(\delta\) of \(a\). So its like you surround \(a\) with an open ball.

OpenStudy (zzr0ck3r):

or circle...

OpenStudy (ikram002p):

They are two important conditions in metric I think we should mention them right ?

OpenStudy (zzr0ck3r):

3, he didnt ask about metrics...

OpenStudy (zzr0ck3r):

The concept stays the same through different metrics but they no longer match the name. The taxi cab metric will gave an open ball that looks like a diamond, and the infinity metric a square.... Also note that the actual ball is not the border, its all the stuff inside(for an open ball). So if you take a basket ball and fill it with air, then the open ball is the air(in R^3 with euclid metric). I hope that makes sense.

OpenStudy (zzr0ck3r):

I do not know what a spare is

OpenStudy (zzr0ck3r):

repost that

OpenStudy (anonymous):

let E =R endowed with the metric Do defined by Do(x,y)= 1, if x is not y and 0 if x =y for arbitrary x,y element of R . compute the ball B(1;1/2)

OpenStudy (zzr0ck3r):

well that says that everything but 1, has a distance of 1 from 1. And we want to know about the points within 1/2 of 1. So there is only one. what is it?

OpenStudy (zzr0ck3r):

Do(1,1.4)=1 Do(1,1.2)=1 Do(1, 0.8)=1 Do(1,1)=0 Do(1,0.6)=1

ganeshie8 (ganeshie8):

lol it is a weird metric, everything is at 1 unit away from everything

OpenStudy (zzr0ck3r):

how many points have a distance of less than 1?

OpenStudy (zzr0ck3r):

not everything :)

ganeshie8 (ganeshie8):

Ah except the self

OpenStudy (zzr0ck3r):

Right, so here is a metric that gives an open ball that is a singleton. Does this make sense? @GIL.ojei ?

OpenStudy (anonymous):

So Sir , what is the question asking us to find and how did u get all those point s and equate them to 1 and how was tour conclusion made??

OpenStudy (zzr0ck3r):

\(B_{Do}(1;1/2)=\{x\in R \mid Do(1,x)<\frac{1}{2}\}=\{1\}\)

OpenStudy (zzr0ck3r):

The question wants the set of all points that are within distance 1/2 of 1. But with this metric, everything, except 1, is distance 1 from 1. So the only point in the set is 1 itself.

OpenStudy (zzr0ck3r):

because the distance from 1 to 1 is 0.

OpenStudy (zzr0ck3r):

if this was the euclidean metric we would have the interval (0.5, 1.5)

ganeshie8 (ganeshie8):

i think "n" points require "n-1" dimensions for this metric to be valid/used

OpenStudy (zzr0ck3r):

I don't know what you mean.

OpenStudy (zzr0ck3r):

It passes all the rules of a metric on a set.

ganeshie8 (ganeshie8):

at least in euclidean metric in \(\mathbb{R}^n\)... if we have 3 points, then they can be at 1 unit away from each other only if they are at corners of an equilateral triangle - two dimensions

ganeshie8 (ganeshie8):

similarly if we have 4 points, we must go to 3-space where the points can be at vertices of a tetrahedron or something .. its hard to visualize for more points idk lol

OpenStudy (zzr0ck3r):

A metric is a binary operation on the set. It takes only two elements as an argument .

OpenStudy (zzr0ck3r):

err not a binary operation but from XxX to R.

ganeshie8 (ganeshie8):

XxX to R is a binary operation which takes two operands as input and spits out one real number as output right

OpenStudy (zzr0ck3r):

I think a binary operation on X has to have X itself as the codomain

OpenStudy (zzr0ck3r):

But yes.

OpenStudy (zzr0ck3r):

\(\circ : X \times X \rightarrow X\)

OpenStudy (zzr0ck3r):

That's a binary operation... but anyway.

OpenStudy (zzr0ck3r):

But I think I see what you are trying to do and that is think about shapes with this metric. I am not willing to take that jump tonight :)

ganeshie8 (ganeshie8):

Exactly! I am trying to visualize, which is forbidden sometimes in real analysis haha!

OpenStudy (zzr0ck3r):

How would we define a square with a normal metric?

ganeshie8 (ganeshie8):

I see your point, taxicab metric works well i think ?

OpenStudy (zzr0ck3r):

I want to think about a square in this metric with that definition.

OpenStudy (zzr0ck3r):

Well not even that. I am just saying how ever we define a square with the normal distance function on R^2, lets use that definition on this metric and try and think of what a square looks like.

ganeshie8 (ganeshie8):

there are only two possible values for distances here : {0, 1}

OpenStudy (zzr0ck3r):

So a square with side length 1, lets say the unit square and look at the point (0,1/2) normally we would have all the points that are 1 unit away in one direction and we get only one point (1,1/2) But with this metric we get EVERYTHING... lol

OpenStudy (zzr0ck3r):

So most shapes will give everything.

OpenStudy (zzr0ck3r):

I think the only purpose of this metric is to ask this question :) ok 5am good night

ganeshie8 (ganeshie8):

Haha that is really weird to visualize! xD

OpenStudy (anonymous):

the metric is just the one that induces the discrete topology, so balls of radius \(r<1\) only contain one point: \(B_{r\,<\,1}(p)=\{p\}\)

OpenStudy (anonymous):

https://en.wikipedia.org/wiki/Discrete_space#Definitions this is because all the points are isolated

OpenStudy (anonymous):

please guys, you have been arguing and i do not understand one bit please, what are the steps in solving the quation i gave and what would be the final answer

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

@oldrin.bataku

OpenStudy (ikram002p):

i need to see the definition in your book of metric space in your book + which class is this to help you more ^_^

OpenStudy (anonymous):

i am in my finals in national open university. here is a link to the book

OpenStudy (ikram002p):

where is it ? :D

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@dan815

OpenStudy (triciaal):

sorry can't help on this I don't know

OpenStudy (anonymous):

let E =R endowed with the metric Do defined by Do(x,y)= 1, if x is not y and 0 if x =y for arbitrary x,y element of R . compute the ball B(1;1/2)

OpenStudy (dan815):

what is the definition of a ball

OpenStudy (anonymous):

the open ball of radius \(r\) about \(p\) in a metric space \((X,d)\) is defined as $$B(p;r)=\{x\in X:d(p,x)<r\}$$

OpenStudy (anonymous):

yes that was the definition in my book, they gave just open balls, closed balls and sphare

OpenStudy (anonymous):

but did not define a ball

OpenStudy (anonymous):

in this case, it doesn't matter whether they ask for open or closed balls, because there are no points other than \(p\) that are up to or within distance \(1/2\) of \(p\), so the ball in our discrete metric is a singleton: $$B(p;1/2)=\{p\}$$

OpenStudy (anonymous):

ok, so what nxt

OpenStudy (anonymous):

remember the definition of the metric here: $$d(x,y)=\left\{\begin{matrix}0&\text{if }x=y\\1&\text{if }x\ne y \end{matrix}\right.$$

OpenStudy (kainui):

@GIL.ojei Explain what you understand.

OpenStudy (anonymous):

yes i remember

OpenStudy (anonymous):

for example, suppose our space consisted of the following points \(p,q,r\). we know: $$d(p,p)=0\\d(p,q)=1\\d(p,r)=1$$ so the only thing within a distance of \(1/2\) is \(p\), since \(d(p,q)=d(p,r)=1>1/2\) and \(d(p,p)=0<1/2\)

OpenStudy (anonymous):

yes, i know that

OpenStudy (anonymous):

okay, and that's it

OpenStudy (anonymous):

that's the problem you asked about, @ganeshie8 answered it hours ago

OpenStudy (anonymous):

and @zzr0ck3r

OpenStudy (anonymous):

so, how did he get does points like d(1;0.8)=1 I MEAN THE 0.8

OpenStudy (anonymous):

OK, WHAT ABOUT THE COMPUTATION OF THIS AND SOLVE COMPLETELY WITH STEPS, PLEASE

OpenStudy (anonymous):

OK, WHAT ABOUT THE COMPUTATION OF THIS AND SOLVE COMPLETELY WITH STEPS, PLEASE

OpenStudy (anonymous):

let E =R endowed with the metric Do defined by Do(x,y)= 1, if x is not y and 0 if x =y for arbitrary x,y element of R . compute the ball B(1;5)

OpenStudy (anonymous):

HELLO

OpenStudy (anonymous):

can some one please answer

OpenStudy (anonymous):

hello

OpenStudy (zzr0ck3r):

You need to tell us when you don't understand us. We were not arguing we were discussing math. What do you not understand?

OpenStudy (anonymous):

does it mean that hat they told us to do is to find points from 1 to <5?

OpenStudy (zzr0ck3r):

Do you understand that \(B(1, \frac{1}{2})\) is the set of all points that are within one half of 1?

OpenStudy (anonymous):

yes that is 1<x<1/2 right?

OpenStudy (zzr0ck3r):

So with the normal metric we would get stuff like 0.6,0.7,0.8,0.9,1.1,1.2,1.3

OpenStudy (zzr0ck3r):

no with the standard metric it would be 0.5<x<1.5

OpenStudy (zzr0ck3r):

Do you see that? \(1\pm0.5\)

OpenStudy (anonymous):

yes, e - neighborhood of 1

OpenStudy (zzr0ck3r):

So this is with the standard metric, but we are not in that metric. In this metric distance works differently than you are used to. The distance between any two different points is 1

OpenStudy (zzr0ck3r):

So now there are no points within 1/2 of 1 (except 1 itself)

OpenStudy (zzr0ck3r):

because everything has distance 1

OpenStudy (anonymous):

So now there are no points within 1/2 of 1 (except 1 itself) ,, please give mare example on it

OpenStudy (zzr0ck3r):

yes exactly

OpenStudy (zzr0ck3r):

the distance between 1 and 3 is 1 the distance between 1 and 7 is 1 the distance between 1 and 900000000000 is 1

OpenStudy (anonymous):

waw

OpenStudy (zzr0ck3r):

So when you ask me what are all the points within 1/2 of 1, I tell you there is only 1 and that is 1 itself

OpenStudy (anonymous):

ok

OpenStudy (zzr0ck3r):

So now you tell me the answer to this question and I will know you understand \(B(56, 0.7)=?\)

OpenStudy (anonymous):

1 following my definition of d(x,y)